OpenStudy (anonymous):

multiply 2x2-10 over x+1 * x+2 over 3x2-15

OpenStudy (anonymous):

then simplify

myininaya (myininaya):

\[\frac{2x^2-10}{x+1}*\frac{x+2}{3x^2-15}=\frac{2(x^2-5)}{x+1}\frac{x+2}{3(x^2-5)}\]

myininaya (myininaya):

the x^2-5 's cancel, right?

myininaya (myininaya):

\[=\frac{2(x+2)}{3(x+1)}\]